### v2

# A Canonical Quantization of Poisson Manifolds: a 2-Groupoid Scheme

Abstract
We canonically quantize a Poisson manifold to a Lie 2-groupoid, complete with a quantization map, and show that it relates geometric and deformation quantization: the perturbative expansion in of the (formal) convolution of two quantized functions yields Kontsevich’s star product. Meanwhile, we can push forward this quantization map (by integrating over homotopies of paths) to obtain a quantization map in traditional geometric quantization. This gives a polarization-free, path integral definition of the quantization map, which does not have a prior prescription for Poisson manifolds and which only has a partial prescription for symplectic manifolds. We show that in conventional quantum mechanics our operators are defined on the entire prequantum Hilbert space and that they all preserve polarized sections, contrary to the Kostant-Souriau approach.
Contents
0 Introduction
Kontsevich proved that every Poisson manifold admits a formal deformation quantization,222Obtained using a sequence of bidifferential operators. called a star product (see [15]). That is, there is an associative product on which deforms the product on in such a way that
This is a perturbative approach to quantization. We can think of the inclusion map
as a perturbative quantization map, which associates to the operator Since this is a perturbative approach we don’t get a -algebra, which is important in physics.
A non-perturbative approach to quantization, which was initiated by Weinstein in [26], is geometric quantization of Poisson manifolds. This quantization method aims to define a -algebra of observables — it is given by polarized sections of a twisted convolution algebra of the symplectic groupoid (the groupoid which integrates the Poisson structure, [11]). In the symplectic case, we would like this -algebra to act on polarized sections of the prequantum line bundle.333For a discussion on when a star product acts on the space of sections of such a quantization, see [18], [19].
Geometric quantization is much less understood than formal deformation quantization (which is consistent with the current understanding of QFT) and is incomplete as it doesn’t prescribe a quantization map.444Here, we mean an injective -linear map from to the -algebra of observables that determines a star product. In this sense, the difficulty of geometric quantization is complementary to the difficulty of formal deformation quantization: one (partially) prescribes a -algebra but not a quantization map, the other prescribes a formal quantization map but not a -algebra.
In some examples (eg. [1],[3],[22]) it has been shown that there exists a quantization map which is compatible with formal deformation quantization, but an explanation of why and when this occurs is lacking. Even in traditional geometric quantization of symplectic manifolds the quantization map is, at best, defined only on a small subspace (and assumes a polarization exists). It is desirable to have a more mathematically precise theory which directly relates these two quantization methods.
The constructions we make are justified by results of the Poisson sigma model in [2], [5]. We formally define a quantization map (definition 3.2.2) by using the Lie 2-groupoid integrating a Poisson manifold ([27]) — this involves computing a path integral over homotopies of algebroid paths. In the symplectic case, this 2-groupoid is a quotient of the singular simplicial set of the manifold. Noncommutativty for arises due to distinct points in phase space being isomorphic in the 2-groupoid. This is a two dimensional sigma model explanation of quantization, somewhat as in [10].
0.1 Main Example
We compute the path integrals in the case of the pointed Kähler manifold with the canonical symplectic form (section 4).555The only examples of geometric quantization of symplectic manifolds that have been shown to have an injective quantization map are discrete quotients of and it is not difficult to compute the path integrals there. The quantization of a smooth function is a section of a trivializable line bundle over the symplectic groupoid, and can be identified with (eq. 4.0.8)
where is the (signed) area of the polygon with vertices and the integration is over
The canonical connection on the prequantum line bundle is (definition 3.3.1) and parallel transport determines a representation of these quantized functions as integral operators on The result is
where and is the (signed) area of the geodesic triangle determined by the integral is over with respect to the product measure.
The prequantum Hilbert space contains both pure and mixed states.666In the -algebra sense. A pure vector state is one which generates an irreducible representation ([9]). Polarized sections form irreducible subrepresentations for all covariantly constant Lagrangian polarizations. That is, if is polarized (with respect to some complex linear polarization), then (eq. 4.0.19)
and this is a polarized section. In particular, we get the position, momentum and Segal-Bargmann (holomorphic) representations. The states of the latter take the form
With the Kostant–Souriau prescription, only a small subspace of functions prequantize to operators which preserve polarized sections, and this subspace depends on the polarization — only affine functions prequantize to operators which preserve all linearly polarized sections. Our operators 0.1.2 are defined polarization-free and they all preserve polarized sections, for all complex linear polarizations. These two quantizations agree on the subspace where the Kostant–Souriau prescription is defined.
In addition, with respect to the twisted convolution algebra on the groupoid and the non-perturbative Moyal product (lemma 1.1.1, eq. 4.0.11), 777The image of under the van Est map is ([17]) and is therefore independent of the complex structure.
Remark 0.1.1.
We view the role of Lagrangian polarizations in the usual definitions of geometric quantization as a tool to help guess a quantization map, but not necessarily as an essential ingredient (sometimes they don’t even exist). This is consistent with Kontsevich’s and Fedosov’s ([8]) constructions of the star product, where polarizations are absent. We will only use them to help construct irreducible representations.
In the usual approach, a common heuristic is that the prequantum Hilbert space being “too large” is a problem. However, our perspective is that the only problem is that the subalgebra of operators generated by the Kostant-Souriau prequantization map is far too large to represent observables — its image only contains first order differential operators. A precise version of this is that the Kostant-Souriau prequantization map is not compatible with a star product (definition 0.3.1). This perspective is particularly useful when the prequantum Hilbert space contains inequivalent irreducible representations.
0.2 Lie 2-Groupoid Approach to Quantization and Main Results
Our 2-groupoid approach to quantization relates formal deformation quantization and geometric quantization of Poisson manifolds, and is closely related to the Poisson sigma model. It has the advantage that it can be defined on a lattice. The idea is based on the following: every Poisson manifold has a canonical integration given by a Lie 2-groupoid, denoted (see [27]). This Lie 2-groupoid comes with a canonical multiplicative prequantum line bundle
Its space of sections, denoted is formally a -algebra — its product is given by convolution (this involves a path integral) and its involution is complex conjugation. It’s norm is (formally) the -norm of the left regular representation.888Every algebra acts on itself on the left, forming a representation called the left regular representation. There is well-defined, canonical quantization map
In this paper we regard a Poisson manifold as a smooth family of classical systems, and therefore we want a corresponding smooth family of quantizations (which often exist formally, but not always. See section 0.3). In this context, we take the pair
to be a “universal” quantization of We have the following theorem:
Remark 0.2.1.
We want to emphasize that our use of the Lie 2-groupoid has nothing to do with the existence of the symplectic groupoid — we need to use the 2-groupoid in order to define
Theorem 0.2.2.
(formal) Let be a Poisson maniold such that the convolution of two sections of is smooth.999This should be true in a wide class of examples, including regular Poisson manifolds satisfying an integrality condition. We have that
1.
For is equal to the 3-point function
(0.2.4) of the Poisson sigma model on the disk ([5]). Therefore, perturbatively at
(0.2.5) -
2.
In the case that the symplectic groupoid is prequantizable, descends to a multiplicative prequantum line bundle
(0.2.6) and pushes forward to a “geometric” quantization map
(0.2.7) -
3.
In the case that is pointed, simply connected and symplectic, there is a canonical representation of on sections of the prequantum line bundle hence there is a canonical representation of the -algebra generated by
To clarify the different (but related) quantization maps:
1.
The map is the 2-groupoid quantization map. The others are obtained from it.
2.
The map in part 2 is the geometric quantization map.
3.
In part 1, the left regular representation defines a third quantization map
(0.2.8) We describe the -algebra which belongs to in the next section, see definition 0.3.3; we denote it by
4.
From the representation in part 3, we get a fourth quantization map by composing with the map into the bounded linear operators on the Hilbert space of the prequantum line bundle.
The main goal is to realize theorem 0.2.2 non-perturbatively, which we are optimistic about in the case of regular Poisson manifolds. This is the goal of the lattice framework recently developed in [17] for defining path integrals over spaces of Lie algebroid morphisms, see appendix 5. Most known examples of non-perturbative deformation quantizations fit into this framework, as discussed in example 9.2.7 of the aforementioned paper. In conventional quantum mechanics, we work out this result in section 4.
In the next section we will make clear exactly what we mean by a quantization.
0.3 Non-Perturbative (Tangential) Deformation Quantizations
We will discuss non-perturbative deformation quantizations. The definition we use is essentially the most general one that is compatible with both -algebraic quantization and formal deformation quantization. It sits between Rieffel’s strict quantization ([20]) and his strict deformation quantization.
Definition 0.3.1.
(see [12]) Let be a Poisson manifold and let be a set containing as an accumulation point. For each let be a unital -algebra such that and let
be injective and - linear such that its image generates Furthermore, assume that is the inclusion map. We say that is a (non-perturbative) deformation quantization of if there is a star product on such that, for all
where is the component of the star product up to order
In this paper we view a Poisson manifold as a smooth family of classical systems, thus we want a quantization which quantizes each of the classical systems. This motivates the next definition:
Definition 0.3.2.
We say that a (non-perturbative) deformation quantization of is tangential if for all whenever are supported on disjoint symplectic leaves.
Every regular Poisson manifold has a tangential, formal deformation quantization,101010A star product is tangential if the sequence of bidifferential operators are tangent to the symplectic leaves. see the work of Fedosov [8], also see [13], [25]. (there are a number of examples of non-regular Poisson manifolds which also have such a quantization, as discussed in [17]).111111A quantization which isn’t tangential exhibits a kind of tunneling. The maps discussed in section 0.2 have a tangential property since the orbits of the Lie 2-groupoid are the symplectic leaves.
A Poisson manifold has a natural -algebra associated to it, which in the case of a symplectic manifold is the space of bounded linear operators on square integrable functions. For a Poisson manifold, it is essentially the direct product of the -algebras of its symplectic leaves:
Definition 0.3.3.
Let be a Poisson manifold. We have a norm on the subspace of measurable functions for which
where the supremum is over all symplectic leaves and is the -norm of pulled back to 121212With respect to the measure induced by the symplectic form. We denote this space by
Let denote the space of bounded linear operators on such that, if for a symplectic leaf
then
There is an involution on defined by
This turns into a -algebra.
If then
Motivated by theorem 0.2.2 and eq. 0.2.8, we are optimistic that regular Poisson manifolds such that the Poisson structure is integral over spheres (ie. satisfies 3.2.1) have non-perturbative, tangential deformation quantizations for which
This would give a a non-perturbative form of Kontsevich’s star product on such Poisson manifolds.131313For examples of non-tangential quantizations, see the quantizations of the dual of nilpotent Lie algebras [21].
Remark 0.3.4.
In formal deformation quantization, it’s really the gauge equivalence class of the formal quantization map that’s important. Similarly, we might define two quantization maps to be gauge equivalent if there is an automorphism of -algebras intertwining them.
1 The Star Product and the Poisson Sigma Model
A Poisson sigma model construction of Kontsevich’s star product as a 3-point function in a QFT was explicitly written out by Cattaneo and Felder in [5].
Precisely, the domain of this Poisson sigma model is the disk with three cyclically ordered marked points on the boundary, denoted The target space is a Poisson manifold and a field is a vector bundle morphism
covering There is a boundary condiition which says that the pullback of to takes values in the zero section. The action is given by
The star product is then given by141414More precisely, we can say that the perturbative expansions at is Kontsevich’s star product.
Here, and the integral is over all fields with The measure is normalized so that
1.1 The Tangential Star Product
Here, we discuss the tangential property of eq. 1.0.2. In the following lemma, is the projection and the domain of integration is the set of Lie algebroid morphisms where the Lie algebroid structure on is determined by the Poisson structure ie. the bracket on exact sections of is
and for the anchor is ([7]).
The action is given by
Lemma 1.1.1.
(see [2]) In the following, if the quantity on the right is smooth on then we have that
The right side of 1.1.2 is exactly what one obtains by restricting the domain of integration of 1.0.2 to Lie algebroid morphisms, ie. classical solutions of the Euler-Lagrange equations. This is tangential because Lie algebroid morphisms stay in a symplectic leaf.
The problem of computing non-perturbatively is to compute (as a distribution on ) the integral kernel
where the domain of integration is the set of Lie algebroid morphisms such that
A lattice approach to defining these path integrals is described in [17], see appendix 5. On a Kähler manifold this kernel should be canonically defined, as there is a canonical 2-cocycle on the local pair groupoid (this is consistent with the fact that Fedosov’s quantization is canonically defined on a Kähler manifold, as the Levi-Civita connection is a symplectic connection). See LABEL:moyal.
2 Geometric Quantization of the Lie 2-Groupoid
Based on eq. 1.1.2, we are going to interpret eq. 1.0.2 using the twisted convolution algebra of a Lie 2-groupoid, obtained via a geometric quantization-like construction. Much of this was explained in [16].
If is a Poisson manifold then is naturally a Lie algebroid, as discussed in the previous section. We can integrate it to a Lie 2-groupoid.
Definition 2.0.1.
We let denote the geometric realization of the standard n-simplex.
We now define the 2-groupoid, see remark 2.0.3 for more details
Definition 2.0.2.
(see [27]) Let be a Poisson manifold. There is a Lie 2-groupoid, denoted which in degree is given by in degree is given by
and in degree is given by
where identifies two morphisms if they are homotopic relative to the boundary. We denote the face maps
by where We denote the face maps
by where
are the 1-dimensional faces of For such an we may write
Remark 2.0.3.
In this definition, means that takes values in the zero section when restricted to the pullback of the tangent bundle to the vertices. Lie algebroid morphisms are homotopic relative to the boundary if there is a Lie algebroid morphism
which agrees with on respectively, and which takes values in the zero section on
A Poisson structure defines a 2-cocycle in the Chevalley-Eilenberg complex of This integrates to a 2-cocycle on the 2-groupoid,151515One should interpret this as the inverse of a higher groupoid van Est map. which defines a multiplicative line bundle
Definition 2.0.4.
We have a 2-cocycle on given by
We then have a multiplicative line bundle
whose multiplication is defined as follows: for composing to via we define
Sections of are denoted
In the above definition, if for all morphisms then the multiplication only depends on That we have defined a cocycle follows from Stokes’ theorem.
Next, we define the twisted convolution algebra:
Definition 2.0.5.
Consider two sections ie. maps Their twisted convolution
is given by161616We are suppressing the dependence of on in the notation.
Explicitly,
This definition is formal and makes sense for any Lie 2-groupoid if we think of as just being a 2-morphism. If we apply it to a Lie 2-groupoid which is just a Lie 1-groupoid then (up to normalization) we recover the usual definition of the twisted convolution algebra (see eg. [11]).
3 Main Result
Here we discuss the three parts of theorem 0.2.2.
3.1 Part 1: The 2-Groupoid Quantization Map
Definition 3.1.1.
The following is part 1 of theorem 0.2.2 and is a restatement of lemma 1.1.1, which we take as a fact for regular Poisson manifolds:
Lemma 3.1.2.
Let be a Poisson manifold for which the right side of 1.1.2 is smooth and let Then (here means the pullback to the space of objects).
3.2 Part 2: The Geometric Quantization Map
If for any morphism we have that
then the multiplicative line bundle of definition 2.0.4 descends to a multiplicative line bundle over the source simply connected, symplectic groupoid (the underlying line bundle was constructed in [2]).
Definition 3.2.1.
Recall that arrows in are equivalence classes of algebroid paths up to homotopy, relative to the endpoints ([7]). We denote such an arrow by Assuming eq. 3.2.1, the following defines a multiplicative line bundle
The vector space over an arrow
is given by equivalence classes of points (defined in definition 2.0.4), such that if then
where is any algebroid homotopy between The multiplication of descends to a multiplication of
We now define the quantization map:
Definition 3.2.2.
Assuming eq. 3.2.1, we define the geometric quantization map by
where the integral is over disks with boundary components ( is fixed). This definition is independent of the representative of
3.3 Part 3: Representation on Sections of Prequantum Line Bundle
On a pointed, simply connected symplectic manifold there is a canonical action of the multiplicative prequantum line bundle of on the prequantum line bundle (with connection)
The construction of the prequantum line bundle is obtained by constructing the line bundle over 171717 is the pair groupoid, it has a unique arrow between any two points in as in the previous section, and pulling it back to the source fiber over We will recall the standard construction in geometric quantization below ([24]).
Assume that Consider the space of paths
Since is simpy connected, the map
descends to a diffeomorphism on homotopy classes of paths relative to the endpoints, The trivial line bundle descends to where we quotient out by
if is a disk whose boundary has components We may therefore denote points in by and points in the line bundle over by
Definition 3.3.1.
The prequantum line bundle is the line bundle defined by the previous construction. The trivial connection on descends to a connection on let and let
be a path such that
Parallel transport over sends
where is any disk whose boundary has components The space of sections is denoted
We are now ready to define the action of on
Definition 3.3.2.
We have an action of on where
acts by parallel transport over 181818We are implicitly using the isomorphism so that these algebroid morphisms into are the same as maps of manifolds into followed by multiplication by This descends to an action of on
Finally, we can discuss the representation:
Definition 3.3.3.
Let We have a representation of on defined by
where
4 Example: Conventional Quantum Mechanics
With the constructions made in the previous section, we are going to explicitly show what theorem 0.2.2 gives in the example of conventional quantum mechanics. We will start with the Kähler structure and determine, in order: the quantization map, the non-perturbative Moyal product and the standard irreducible representations. The path integrals can be computed using the framework developed in [17], and we will assume the results (see appendix 5 for the idea).
Consider the pointed Kähler manifold 191919All of our constructions will depend only on this pointed Kähler structure.
1.
The symplectic groupoid is
(4.0.1) We are identifying the fundamental groupoid of a simply connected space with its pair groupoid (which has a unique arrow between any two points).
2.
The multiplicative line bundle in definition 3.2.1 is isomorphic to the trivial complex line bundle twisted by the 2-cocycle given by
(4.0.2) which is the (signed) area of the geodesic triangle with vertices This is the natural -cocycle obtained using the Kähler structure.202020The van Est map applied to gives The multiplication of the line bundle is given by
(4.0.3) -
3.
There is a natural non-vanishing section of the prequantum line bundle in definition 3.3.1 obtained using the Kähler structure. It is given by
(4.0.4) where we identify with the geodesic connecting to This determines a trivialization of which we will assume from now on. In this trivialization, the canonical connection is given by
(4.0.5) We have implicitly used the “geodesic” trivialization of in step 2.
4.
The action of on in definition 3.3.2, obtained using parallel transport, is given by
(4.0.6) We have that
5.
The representation of the corresponding twisted convolution algebra in definition 3.3.3 on is given by
(4.0.7) -
6.
The quantization map in definition 3.2.2 is given by212121We have normalized by dividing out by eq. 1.1.3.
(4.0.8) Geometrically, the exponent is equal to the (signed) area of the polygon determined by the points
(4.0.9) Thus, equivariant under automorphisms (of the pointed Kähler structure).
7.
Note that there is a Lagrangian polarization induced by the morphism
(4.0.10) and is polarized with respect to the connection 222222Compare this with the construction of the quantization in [11], for which a polarization needs to be chosen. The two quantization maps are related by an automorphism of
8.
The product (defined in lemma 1.1.1) is given by
(4.0.11) (4.0.12) This is equal to in theorem 0.2.2 and is the non-perturbative Moyal product ([22], [28]). It is invariant under automorphisms of the Kähler manifold and the integral kernel is which is equal the integral kernel in eq. 1.1.3.
9.
We have that and in fact
This completes the constructions of theorem 0.2.2. We will now discuss irreducible subrepresentations of our -algebra, using eq. 4.0.13.
10.
Covariantly constant Lagrangian polarizations are of the form
(4.0.14) for constants (not both ). A polarized section can be written as
for any such that
(4.0.16) With respect to the Kähler polarization, polarized sections are of the form
These give irreducible subrepresentations of the left representation of ie. is of the form 4.0.15, for the same and for any See eq. 4.0.19. -
11.
The actions of the quantizations of on (as in 4.0.15) are given by242424Our conventions imply that
(4.0.17) (4.0.18) We can read off the position, momentum and Segal-Bargmann representations (ie. the holomorphic representation [14]).252525The holomorphic representation is the only one which forms a -algebra subrepresentation, the former are only algebra subrepresentations, but there is essentially a unique measure turning them into -algebra representations.
We are describing the representation of the quantum observables on all vector states262626A pure state is one which generates an irreducible representation, these polarized sections are pure states. However, the quantization map is defined on all sections of the prequantum line bundle, and some of these sections are mixed states. in the sense of -algebras. See [9]. before choosing a polarization. In traditional quantization, one chooses a polarization first and the Kostant-Souriau quantization prescription only quantizes a small subspace of observables which preserves polarized sections.
Generally, if is a polarized section, then(4.0.19) -
12.
The involution lifts to an isomorphism of representations from states polarized with respect to to states polarized with respect to The isomorphism is given simply by integration with respect to where we identify the vector spaces over and using the flat connection along the polarization, ie. under parallel transport over we have that
(4.0.20) Performing the integral, we get
4.1 The Classical Limit
We can analyze the classical limit from the 2-groupoid perspective. First, note that given a Poisson structure on we get a natural Poisson structure on where the induced Poisson structure on is .272727We can replace with any subset with as an accumulation point. Then, what we have really done in this paper is integrated the extension of Lie algebroids282828 is the trivial line bundle over
to a multiplicative line bundle
This extension is the one obtained by interpreting as a 2-cocycle on
The case is special: isomorphism classes (ie. orbits) of objects in the 2-groupoid of are its symplectic leaves, and is the unique case where the symplectic leaves are just the points in As a result, the pure states at are just pointwise evaluation. From the 2-groupoid perspective, the measurement problem is due to the existence of isomorphisms between distinct points in phase space of a symplectic manifold, for
Even though algebroid paths are constant at there is still nontrivial dynamics since we aren’t varying the 2-cocycle with thus the case still remembers the Poisson structure. This is the only case where the -cocycle isn’t proportioinal to the canonical 2-cocycle on (since when the canonical 2-cocycle is ).
One can take a line bundle over which prequantizes each of its symplectic leaves. A representation of the quantization on this space of sections will pull back to a representation on sections of the line bundle over each symplectic leaf (a “subrepresentation”). It is natural to take a subspace of sections over which pull back to irreducible representations over each leaf (eg. by taking a foliation by coisotropic submanifolds whose intersection with each symplectic leaf is Lagrangian). We can think of this as a classical mixture of quantum systems.
5 Concluding Remarks
In our main example, one can check that if is periodic and transforms as a theta function under translations, then transforms as a theta function. Therefore, if we carry out our construction on the symplectic torus, we will get a representation of the noncommuative torus on sections of the prequantum line bundle. It would be interesting to construct the irreducible representations.
In [23] (see also [4], [6]) it is shown that there is a sequence of “quantization maps” and representations associated to a compact Kähler manifold (Berezin-Toeplitz quantization) which induce a formal deformation quantization. However, these quantization maps are far from injective. We expect that our construction can be carried out in this context, and that their representations factor through our geometric quantization map with the path integrals being computed as in [17].
We will briefly describe the definition of path integrals over spaces of Lie algebroid morphisms given in [17], specialized to the example of the star product in the form eq. 1.1.2, ie.
The steps are:
1.
triangulate the disk (with the marked points contained in the vertices) and form the associated simplicial set
2.
integrate the Poisson manifold to the local symplectic groupoid,
3.
approximate the space of Lie algebroid morphisms by morphisms of simplicial sets between and the local symplectic groupoid,
4.
integrate the Poisson structure to a 2-cocycle, via the van Est map,
5.
form the approximations of the action (generalized Riemann sums),
6.
construct the measure on the hom space of simplicial sets by using the symplectic form on the symplectic leaves and the Haar measures on the isotropy groups,
7.
compute the approximations of the functional integral,
8.
take a limit over triangulations of the disk.
The path integrals in 4 can then be computed using the integral representation of the delta function.